OpenStudy (anonymous):

If a rock falls from a height of 90 meters on Earth, the height H (in meters) after x seconds is approximately H(x)=90-4.9x^2. When does the rock strike the ground? Can someone explain the process of getting the answer to me?

OpenStudy (stormfire1):

The formula in this case is distance = 1/2 g t^2. You know the distance is 90m and g = 9.8 m/s so you can just solve it algebraically from there. I'm getting ~4.28 seconds

OpenStudy (stormfire1):

To be perfectly clear: 2(90/9.8) = t^2...then just take the square root of that result to get the time in seconds.